### v1

# Weak Coupling and Spectral Instability for Neumann Laplacians

Abstract.
We prove an abstract criterion on spectral instability of nonnegative selfadjoint extensions of a symmetric operator and apply this to self-adjoint Neumann Laplacians on bounded Lipschitz domains, intervals, and graphs. Our results can be viewed as variants of the classical weak coupling phenomenon for Schrödinger operators in $L^{2}({\mathbb{R}}^{n})$ for $n=1,2$.
Key words and phrases:
Weakly coupled bound state, spectral instability, Birman–Schwinger principle, Neumann Laplacian, Schrödinger operator
2020 Mathematics Subject Classification:
Primary: 35J10, 47A10, 47A75, 47F05, 81Q10; Secondary: 34L05, 35P05, 47A55, 47E05.
1. Introduction
Let us start by recalling the classical weak coupling phenomenon for Schrödinger operators, which goes back to Simon [30], [31]. For this purpose, let $-\Delta$ be the self-adjoint one-dimensional Laplacian in $L^{2}({\mathbb{R}})$ defined on $H^{2}({\mathbb{R}})$ and assume that the potential $V:{\mathbb{R}}\rightarrow{\mathbb{R}}$ satisfies $V\in L^{1}({\mathbb{R}};(1+x^{2})dx)\cap L^{2}({\mathbb{R}};dx)$, and $V$ is not zero a.e. For $\alpha\in{\mathbb{R}}$ it follows that $-\Delta+\alpha V$ is self-adjoint in $L^{2}({\mathbb{R}})$ and
$\sigma_{ess}(-\Delta+\alpha V)=\sigma_{ess}(-\Delta)=\sigma(-\Delta)=[0,\infty).$ (1.1)
It was shown in [30, Theorem 2.5] (see also [29, Theorem XIII.11]) that for any $\alpha<0$ one has
[…]
and hence, in particular, if $V\geq 0$, then $\sigma_{p}(-\Delta+\alpha V)\cap(-\infty,0)\not=\emptyset$ for any $\alpha<0$. The same result holds also for the self-adjoint Laplacian $-\Delta$ in $L^{2}({\mathbb{R}}^{2})$ under slightly different integrability conditions on the potential $V:{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$, and it is also well known that the phenomenon of weakly coupled eigenvalues does not appear in dimensions $n\geq 3$. The works [30], [31] by Simon have inspired and influenced a lot of future research; they were followed by Klaus and Simon [21], [22], and Rauch [28]. A wealth of additional information can be found, for instance, in [5], [6], [8], [13], [18][20], [25], [26], [27], [29, Theorem XIII.11, p. 336–338].
The main objective of this note is to transfer these ideas from Schrödinger operators $-\Delta+\alpha V$ to an abstract setting that replaces the Laplacian by a nonnegative self-adjoint extension $A$ of a densely defined closed symmetric operator $S\geq\kappa I_{{\mathfrak{H}}}$, $\kappa>0$, in a Hilbert space ${\mathfrak{H}}$ and the potential by an appropriate nonnegative self-adjoint perturbation, also denoted by $V$, that is relatively form compact with respect to $A$. In our main abstract result Theorem 2.2 it is shown that under some additional mild assumptions $A$ is spectrally unstable, that is, for any $\alpha<0$ the perturbed self-adjoint operator $A+\alpha V$ has negative discrete eigenvalues. The proof of Theorem 2.2 is based on the Birman–Schwinger principle, see, for instance, [9, 23]. In fact, the essential assumptions to ensure the existence of weakly coupled negative eigenvalues of $A+\alpha V$, $\alpha<0$, are $\ker(A)\not=\{0\}$ and $\ker(A)\not\subseteq\ker(V)$; roughly speaking the first assumption $\ker(A)\not=\{0\}$ ensures that the resolvent of $A$ has a singularity at 00 and the second assumption $\ker(A)\not\subseteq\ker(V)$ is needed to preserve this singularity for the sandwiched resolvent $V^{1/2}(A-\mu I_{{\mathfrak{H}}})^{-1}V^{1/2}$ when $\mu<0$ tends to 00. We note that for the special case where 00 is an isolated eigenvalue of finite multiplicity of $A$, our result would also follow from asymptotic perturbation theory.
Our general result applies directly to the Neumann Laplacian $-\Delta_{N}$ on a bounded interval $(a,b)$ or on a bounded Lipschitz domain $\Omega\subset{\mathbb{R}}^{n}$, $n\geq 2$, since in that case $\ker(-\Delta_{N})$ is spanned by the constant function and if $V\geq 0$ is a multiplication operator which is relatively compact perturbation with respect to $-\Delta_{N}$, then certainly $\ker(-\Delta_{N})\not\subseteq\ker(V)$ as otherwise $V=0$ a.e. on $(a,b)$ or $\Omega$, respectively. As a consequence of our abstract result, Theorem 2.2, we conclude in Corollary 3.1 and Corollary 3.2 that for any $\alpha<0$ and nonnegative function $V$, $V\not=0$, such that $V\in L^{p}$ with $p\geq 2$ if $n=1,2$ and $p>2n/3$ if $n\geq 3$, there exist weakly coupled negative bound states for the perturbed Neumann Laplacian $-\Delta_{N}+\alpha V$, that is,
$\sigma(-\Delta_{N}+\alpha V)\cap(-\infty,0)\not=\emptyset\quad\text{for any}\,\,\alpha<0.$ (1.3)
We note that weakly coupled bound states for Schrödinger operators in ${\mathbb{R}}^{n}$ exist only for $n=1,2$, whereas weakly coupled bound states for the perturbed Neumann Laplacian exist in any space dimension. We mention that our abstract result also applies to other self-adjoint nonnegative realizations $A$ of the Laplacian on bounded domains with the property $\ker(A)\not=\{0\}$ (cf. Remark 3.3). The observations for the case of a bounded interval extend naturally to finite compact graphs, where the Neumann Laplacian corresponds to Kirchhoff or standard boundary conditions; see, Corollary 3.4. Furthermore, in Corollary 3.5 we consider a Sturm–Liouville operator with Neumann boundary conditions in $L^{2}((0,\infty))$ with 00 as embedded eigenvalue at the bottom of the essential spectrum.
Finally, a few remarks about the notation employed: Given a separable complex Hilbert space ${\mathfrak{H}}$, $(\,\cdot\,,\,\cdot\,)_{{\mathfrak{H}}}$ denotes the scalar product in ${\mathfrak{H}}$ (linear in the second factor), $\|\,\cdot\,\|_{{\mathfrak{H}}}$ the norm in ${\mathfrak{H}}$, and $I_{{\mathfrak{H}}}$ represents the identity operator on ${\mathfrak{H}}$. The domain and range of a linear operator $T$ in ${\mathfrak{H}}$ are abbreviated by dom (${\text{\rm dom\,}}(T)$ and ran (${\text{\rm ran\,}}(T)$. The kernel (null space) of $T$ is denoted by $\ker(T)$. The spectrum, point spectrum (i.e., the set of eigenvalues), essential spectrum, and resolvent set of a self-adjoint operator in ${\mathfrak{H}}$ will be abbreviated by $\sigma(\,\cdot\,)$, $\sigma_{p}(\,\cdot\,)$, $\sigma_{ess}(\,\cdot\,)$, and $\rho(\,\cdot\,)$, respectively. The space of compact linear operators in ${\mathfrak{H}}$ is denoted by ${\mathcal{B}}_{\infty}({\mathfrak{H}})$. For $\Omega\subseteq\mathbb{R}^{n}$, $n\in\mathbb{N}$, we will abbreviate $L^{2}(\Omega;d^{n}x)$ for simplicity by $L^{2}(\Omega)$, and $I_{L^{2}(\Omega)}$ for convenience by $I$.
2. Spectral instability of nonnegative self-adjoint extensions
Throughout this section suppose that $S$ is a densely defined closed symmetric operator in a Hilbert space ${\mathfrak{H}}$ and assume that $S$ is semibounded from below with the lower bound $\kappa\geq 0$, that is,
$(Sf,f)_{{\mathfrak{H}}}\geq\kappa(f,f)_{{\mathfrak{H}}},\quad f\in{\text{\rm dom\,}}(S).$ (2.1)
Hypothesis 2.1.
Let $A$ be a nonnegative self-adjoint extension of $S$ in ${\mathfrak{H}}$ such that $\ker A\not=\{0\}$ and let $V\geq 0$ be a self-adjoint operator in ${\mathfrak{H}}$ which is relatively compact with respect to $A$, that is,
${\text{\rm dom\,}}(A)\subseteq{\text{\rm dom\,}}(V)\quad\text{and}\quad V(A+I_{{\mathfrak{H}}})^{-1}\in{\mathcal{B}}_{\infty}({\mathfrak{H}}).$ (2.2)
One notes that the lower bound of $A$ is $\kappa\geq 0$ and recalls that the Friedrichs extension $A_{F}$ of $S$ has the same lower bound as $S$. In the case of differential operators (see the next section) the reader may think of $A$ in Hypothesis 2.1 as the self-adjoint Laplacian with Neumann boundary conditions. Another typical example for a self-adjoint extension of $S$ satisfying Hypothesis 2.1 is the Krein–von Neumann extension $A_{K}$, the smallest nonnegative extension of $S$, which in the case $\kappa>0$ is defined by
$A_{K}=S^{*}\upharpoonright{\text{\rm dom\,}}(A_{K}),\quad{\text{\rm dom\,}}(A_{K})={\text{\rm dom\,}}(S)\overset{\textbf{\Large.}}{+}\ker(S^{*})$ (2.3)
(see, e.g., [2, Sect. 5.4], [7] and the references cited therein). We will return to this topic elsewhere.
Our goal is to show that the lower bound 00 for $A$ in Hypothesis 2.1 is not stable under arbitrary small negative perturbations $\alpha V$. The relative compactness assumption in Hypothesis 2.1 ensures that the operators $A+\alpha V$, $\alpha\in{\mathbb{R}}$, are self-adjoint and that
$\sigma_{ess}(A+\alpha V)=\sigma_{ess}(A),$ (2.4)
see, for instance, [29, Theorem XIII.14 and Corollary 2]. Furthermore,
  • $(i)$
    If $\alpha\geq 0$, then $A+\alpha V\geq 0$ and, in particular, $\sigma(A+\alpha V)\cap(-\infty,0)=\emptyset$.
  • $(ii)$
    If $\alpha<0$, then $\sigma(A+\alpha V)\cap(-\infty,0)$ is either empty or consists of discrete eigenvalues.
From Hypothesis 2.1 one obtains $V(A-zI_{{\mathfrak{H}}})^{-1}\in{\mathcal{B}}_{\infty}({\mathfrak{H}})$, $z\in\rho(A)$, by using the resolvent identity. We also note that
$V^{1/2}(A+I_{{\mathfrak{H}}})^{-1/2}\in{\mathcal{B}}_{\infty}({\mathfrak{H}})$ (2.5)
by [10, Theorem 3.5 $(i)$]. Then one has $V^{1/2}(A+I_{{\mathfrak{H}}})^{-1}\in{\mathcal{B}}_{\infty}({\mathfrak{H}})$ and the resolvent identity shows
$V^{1/2}(A-zI_{{\mathfrak{H}}})^{-1}\in{\mathcal{B}}_{\infty}({\mathfrak{H}})\quad\text{and}\quad V^{1/2}(A-zI_{{\mathfrak{H}}})^{-1/2}\in{\mathcal{B}}_{\infty}({\mathfrak{H}}),\quad z\in\rho(A).$ (2.6)
It follows that $(A-zI_{{\mathfrak{H}}})^{-1}V^{1/2}$ and $(A-zI_{{\mathfrak{H}}})^{-1/2}V^{1/2}$, $z\in\rho(A)$, are densely defined bounded operators, whose closures coincide with the adjoints of the operators in (2.6) for $\bar{z}\in\rho(A)$, and hence also belong to ${\mathcal{B}}_{\infty}({\mathfrak{H}})$. Therefore, the Birman-Schwinger family $K(z)$, defined by
$K(z):=\overline{V^{1/2}(A-zI_{{\mathfrak{H}}})^{-1}V^{1/2}},\quad z\in\rho(A),$ (2.7)
satisfies
$K(z)=V^{1/2}(A-zI_{{\mathfrak{H}}})^{-1/2}\overline{(A-zI_{{\mathfrak{H}}})^{-1/2}V^{1/2}}\in{\mathcal{B}}_{\infty}({\mathfrak{H}}),\quad z\in\rho(A).$ (2.8)
Thus, if $z\in\rho(A)$ and $\alpha^{-1}$ is not an eigenvalue of the compact operator $K(z)$, then $K(z)+\alpha^{-1}I_{{\mathfrak{H}}}$ is boundedly invertible and one verifies in the same way as in [9, Proof of Theorem 2.3] that in the present case $z\in\rho(A+\alpha V)$ and the resolvent formula
$(A+\alpha V-zI_{{\mathfrak{H}}})^{-1}$ $=(A-zI_{{\mathfrak{H}}})^{-1}$ (2.9)
$\quad-\overline{(A-zI_{{\mathfrak{H}}})^{-1}V^{1/2}}\bigl{[}K(z)+\alpha^{-1}I_{{\mathfrak{H}}}\bigr{]}^{-1}V^{1/2}(A-zI_{{\mathfrak{H}}})^{-1},$
$\hskip 156.49014ptz\in\rho(A+\alpha V)\cap\rho(A),$
holds.
The next theorem is our main abstract result; it provides a sufficient condition for spectral instability of the self-adjoint operator $A$ in Hypothesis 2.1.
Theorem 2.2.
Let $A$ and $V$ be as in Hypothesis 2.1 and assume, in addition, that $\ker(A)\not\subseteq\ker(V)$. Then
$\sigma(A+\alpha V)\cap(-\infty,0)\not=\emptyset\,\text{ for any $\alpha<0$.}$ (2.10)
Proof.
By assumption there exists $k\in\ker(A)\subseteq\ker(S^{*})$, $\|k\|_{{\mathfrak{H}}}=1$, such that $Vk\not=0$ and hence also $V^{1/2}k\not=0$. From this we conclude that there exists $f\in{\text{\rm dom\,}}\big{(}V^{1/2}\big{)}$ such that $\big{(}V^{1/2}f,k\big{)}_{{\mathfrak{H}}}\not=0$ as otherwise $k\in\big{(}{\text{\rm ran\,}}\big{(}V^{1/2}\big{)}\big{)}^{\bot}=\ker\big{(}V^{1/2}\big{)}=\{0\}$. If $P$ denotes the orthogonal projection in ${\mathfrak{H}}$ onto $\ker(S^{*})$ and we write $\ker(S^{*})=\textrm{lin.span}\{k\}\oplus{\mathfrak{K}}$, then $P=(\,\cdot\,,k)_{{\mathfrak{H}}}k+P_{\mathfrak{K}}$, where $P_{\mathfrak{K}}$ is the orthogonal projection in ${\mathfrak{H}}$ onto ${\mathfrak{K}}$. Now it follows from $\big{(}V^{1/2}f,k\big{)}_{{\mathfrak{H}}}\neq 0$ that
$h_{0}:=PV^{1/2}f=\big{(}V^{1/2}f,k\big{)}_{{\mathfrak{H}}}k+P_{\mathfrak{K}}V^{1/2}f\not=0.$ (2.11)
We shall now make use of the orthogonal direct sum decomposition
${\mathfrak{H}}={\text{\rm ran\,}}(S)\oplus\ker(S^{*})$ (2.12)
and decompose $h=V^{1/2}f=Sg_{S}+h_{0}$ for some $g_{S}\in{\text{\rm dom\,}}(S)$ and $h_{0}\in\ker(S^{*})$ as in (2.11). For $\nu<0$ it follows from $(A-\nu I_{{\mathfrak{H}}})h_{0}=-\nu h_{0}$ and (2.12) that
$\begin{split}\bigl{(}V^{1/2}(A-\nu I_{{\mathfrak{H}}})^{-1}V^{1/2}f,f\bigr{)}_{{\mathfrak{H}}}&=\bigl{(}(A-\nu I_{{\mathfrak{H}}})^{-1}h,h\bigr{)}_{{\mathfrak{H}}}\\ &=\bigl{(}(A-\nu I_{{\mathfrak{H}}})^{-1}(Sg_{S}+h_{0}),Sg_{S}+h_{0}\bigr{)}_{{\mathfrak{H}}}\\ &=\bigl{(}(A-\nu I_{{\mathfrak{H}}})^{-1}Sg_{S},Sg_{S}\bigr{)}_{{\mathfrak{H}}}-\frac{1}{\nu}(h_{0},h_{0})_{{\mathfrak{H}}}\\ &=\int_{0}^{\infty}\frac{\lambda^{2}}{\lambda-\nu}d(E_{A}(\lambda)g_{S},g_{S})_{{\mathfrak{H}}}-\frac{1}{\nu}\|h_{0}\|^{2}_{{\mathfrak{H}}},\end{split}$ (2.13)
where $E_{A}(\lambda)$, $\lambda\in{\mathbb{R}}$, denotes the family of spectral projections of the self-adjoint extension $A$. It is not difficult to verify that the spectral integral converges for $\nu\uparrow 0$ and as $h_{0}\not=0$ the second term tends to +$\infty$+ ∞ for $\nu\uparrow 0$.
We note that for $\nu<0$ the Birman–Schwinger operator $K(\nu)$ in (2.7)-(2.8) is nonnegative and compact. Furthermore, from (2.13) we conclude that
$\lim_{\nu\uparrow 0}\,\bigl{\|}K(\nu)\bigr{\|}_{{\mathcal{B}}({\mathfrak{H}})}=+\infty$ (2.14)
and since the operator norm of the nonnegative compact operator $K(\nu)$, $\nu<0$, coincides with its largest eigenvalue we conclude that for any $\alpha<0$ there exist $\nu_{\alpha}<0$ and $k_{\alpha}\in{\mathfrak{H}}$, $k_{\alpha}\not=0$, such that
$K(\nu_{\alpha})k_{\alpha}=-\frac{1}{\alpha}k_{\alpha}.$ (2.15)
Now consider $f_{\alpha}=\overline{(A-\nu_{\alpha}I_{{\mathfrak{H}}})^{-1}V^{1/2}}k_{\alpha}$ (see also [23] or [9, Proof of Theorem 3.2] for the following arguments) and observe first that
$\begin{split}k_{\alpha}&=\bigl{[}K(z)+\alpha^{-1}I_{{\mathfrak{H}}}\bigr{]}^{-1}\bigl{[}K(z)-K(\nu_{\alpha})\bigr{]}k_{\alpha}\\ &=(z-\nu_{\alpha})\bigl{[}K(z)+\alpha^{-1}I_{{\mathfrak{H}}}\bigr{]}^{-1}V^{1/2}(A-zI_{{\mathfrak{H}}})^{-1}\overline{(A-\nu_{\alpha}I_{{\mathfrak{H}}})^{-1}V^{1/2}}k_{\alpha}\\ &=(z-\nu_{\alpha})\bigl{[}K(z)+\alpha^{-1}I_{{\mathfrak{H}}}\bigr{]}^{-1}V^{1/2}(A-zI_{{\mathfrak{H}}})^{-1}f_{\alpha},\quad z\in\rho(A),\end{split}$ (2.16)
and hence, in particular, $f_{\alpha}\not=0$ as otherwise $k_{\alpha}=0$. Using (2.16) we see on the one hand
$\begin{split}&\overline{(A-zI_{{\mathfrak{H}}})^{-1}V^{1/2}}k_{\alpha}\\ &\quad=(z-\nu_{\alpha})\overline{(A-zI_{{\mathfrak{H}}})^{-1}V^{1/2}}\bigl{[}K(z)+\alpha^{-1}I_{{\mathfrak{H}}}\bigr{]}^{-1}V^{1/2}(A-zI_{{\mathfrak{H}}})^{-1}f_{\alpha}\\ &\quad=(z-\nu_{\alpha})\bigl{[}(A-zI_{{\mathfrak{H}}})^{-1}-(A+\alpha V-zI_{{\mathfrak{H}}})^{-1}\bigr{]}f_{\alpha},\end{split}$ (2.17)
where (2.9) was used in the last equality. On the other hand, by the resolvent identity one obtains
$\begin{split}&\overline{(A-zI_{{\mathfrak{H}}})^{-1}V^{1/2}}k_{\alpha}\\ &\quad=\overline{(A-\nu_{\alpha}I_{{\mathfrak{H}}})^{-1}V^{1/2}}k_{\alpha}+(z-\nu_{\alpha})(A-zI_{{\mathfrak{H}}})^{-1}\overline{(A-\nu_{\alpha}I_{{\mathfrak{H}}})^{-1}V^{1/2}}k_{\alpha}\\ &\quad=f_{\alpha}+(z-\nu_{\alpha})(A-zI_{{\mathfrak{H}}})^{-1}f_{\alpha}.\end{split}$ (2.18)
It follows from (2.17) and (2.18) that $(\nu_{\alpha}-z)(A+\alpha V-zI_{{\mathfrak{H}}})^{-1}f_{\alpha}=f_{\alpha}$ which implies $f_{\alpha}\in{\text{\rm dom\,}}(A+\alpha V)$ and $(A+\alpha V)f_{\alpha}=\nu_{\alpha}f_{\alpha}$. Hence $\nu_{\alpha}$ is an eigenvalue of $A+\alpha V$, thus $\sigma(A+\alpha V)\cap(-\infty,0)\not=\emptyset$ for any $\alpha<0$. ∎
Remark 2.3.
We note that for the unperturbed nonnegative self-adjoint operator $A$ in Hypothesis 2.1 it is only assumed that $0\in\sigma_{p}(A)$, but no further restrictions on the spectrum of $A$ are required; for example, in general 00 may be an eigenvalue of infinite multiplicity or an accumulation point of positive spectrum of $A$. In the special case where 00 is an isolated eigenvalue of finite multiplicity of $A$, the spectral instability of $A$ in Theorem 2.2 would already follow from well-known results in asymptotic perturbation theory, see, for instance, [16, Theorem 8] or [17, Theorem VIII.4.9].
3. Spectral instability of the Neumann Laplacian
In this section we shall show that Theorem 2.2 applies to the Neumann Laplacian on bounded Lipschitz domains, (arbitrary) intervals, and graphs, and conclude spectral instability for certain classes of potentials $V$ that are relatively compact.
In the following let $\Omega\subset{\mathbb{R}}^{n}$, $n\geq 2$, be a bounded Lipschitz domain and let $\nu$ be the unit normal vector field pointing outwards on $\partial\Omega$. We shall use the notation
$H^{3/2}_{\Delta}(\Omega)=\bigl{\{}f\in H^{3/2}(\Omega)\,\big{|}\,\Delta f\in L^{2}(\Omega)\bigr{\}},$ (3.1)
where $H^{3/2}(\Omega)$ is the $L^{2}$-based Sobolev space on $\Omega$ of fractional order $3/2$. We recall from [1, 12] that the Dirichlet trace mapping $C^{\infty}(\overline{\Omega})\ni f\mapsto f|_{\partial\Omega}$ and the Neumann trace mapping $C^{\infty}(\overline{\Omega})\ni f\mapsto\nu\cdot\nabla f|_{\partial\Omega}$ extend by continuity to continuous surjective mappings
$\tau_{D}:H^{3/2}_{\Delta}(\Omega)\rightarrow H^{1}(\partial\Omega)\quad\text{and}\quad\tau_{N}:H^{3/2}_{\Delta}(\Omega)\rightarrow L^{2}(\partial\Omega),$ (3.2)
respectively, where $H^{1}(\partial\Omega)$ denotes the first-order $L^{2}$-based Sobolev space on $\partial\Omega$. In the next corollary we study the weak coupling behaviour of the Neumann Laplacian
$A_{N}f=-\Delta f,\quad{\text{\rm dom\,}}(A_{N})=\bigl{\{}f\in H^{3/2}_{\Delta}(\Omega)\,\big{|}\,\tau_{N}f=0\bigr{\}},$ (3.3)
which is self-adjoint in $L^{2}(\Omega)$, see, for instance, [1, Theorem 6.10] or [11, Theorem 2.6 and Lemma 4.8] and also [14].
Corollary 3.1.
Let $\Omega\subset{\mathbb{R}}^{n}$ be a bounded Lipschitz domain, $n\in\mathbb{N}$, $n\geq 2$, suppose that $A_{N}$ is the self-adjoint Neumann Laplacian in $L^{2}(\Omega)$, and assume that $V\not=0$ is a nonnegative function such that $V\in L^{p}(\Omega)$ with $p\geq 2$ if $n=2$ and $p>2n/3$ if $n\geq 3$. Then
$(A_{N}+\alpha V)f=-\Delta f+\alpha Vf,\quad{\text{\rm dom\,}}(A_{N}+\alpha V)={\text{\rm dom\,}}(A_{N}),$ (3.4)
is self-adjoint in $L^{2}(\Omega)$,
$V(A_{N}-zI)^{-1}\in{\mathcal{B}}_{\infty}\big{(}L^{2}(\Omega)\big{)},\quad z\in\rho(A_{N}),$ (3.5)
and
$\sigma(A_{N}+\alpha V)\cap(-\infty,0)\not=\emptyset\,\text{ for any $\alpha<0$.}$ (3.6)
Proof.
Consider the densely defined closed symmetric operator
$Sf=-\Delta,\quad{\text{\rm dom\,}}(S)=H^{2}_{0}(\Omega)=\overline{C_{0}^{\infty}(\Omega)}^{\|\cdot\|_{H^{2}(\Omega)}},$ (3.7)
in $L^{2}(\Omega)$ and note that $S$ is semibounded from below by $\kappa>0$, where $\kappa$ is the smallest eigenvalue of the Friedrichs (or Dirichlet) extension
$A_{F}f=-\Delta f,\quad{\text{\rm dom\,}}(A_{F})=\big{\{}f\in H^{3/2}_{\Delta}(\Omega)\,\big{|}\,\tau_{D}f=0\big{\}};$ (3.8)
[…]
It remains to show that $V$ is relatively compact with respect to $A_{N}$ as then Hypothesis 2.1 is satisfied and the statement follows from Theorem 2.2. In order to see that $V$ is relatively compact with respect to $A_{N}$ we shall use that for $10<\delta<$ 1 one has
$\|f\|_{L^{2q}(\Omega)}\leq C_{q}\|f\|_{H^{3/2-\delta/2}(\Omega)}\,\text{ for }\,q\in\begin{cases}[1,\infty]&\text{if}\,\,n=2,\\ [1,n/(n-3+\delta)]&\text{if}\,\,n\in\mathbb{N},\,n\geq 3,\end{cases}$ (3.9)
by [4, Theorem 8.12.6.I]. Let us consider the case $n\geq 3$ first. As $\Omega$ is bounded we have $L^{p_{2}}(\Omega)\subseteq L^{p_{1}}(\Omega)$, $1\leq p_{1}\leq p_{2}\leq\infty$, and hence under our assumptions there exists $10<\delta<$ 1 such that $V\in L^{p}(\Omega)$, where $p=2n/(3-\delta)$. This yields $V\in L^{2r}(\Omega)$, where $r=n/(3-\delta)$. For $s=n/(n-3+\delta)$ we have $11/r+1/s=$ 1 and the Hölder inequality together with (3.9) leads to
$\|Vf\|_{L^{2}(\Omega)}\leq\|V\|_{L^{2r}(\Omega)}\|f\|_{L^{2s}(\Omega)}\leq C_{s}\|V\|_{L^{2r}(\Omega)}\|f\|_{H^{3/2-\delta/2}(\Omega)},$ (3.10)
so that
$V:H^{3/2-\delta/2}(\Omega)\rightarrow L^{2}(\Omega)$ (3.11)
is bounded. In the case $n=2$ it follows in the same way with $V\in L^{2r}(\Omega)$, $r=1$, and $s=\infty$ that the mapping $V$ in (3.11) is bounded.
Next, one observes that $(A_{N}+I)^{-1}:L^{2}(\Omega)\rightarrow H^{3/2}(\Omega)$ is bounded; this follows, for instance, from the norm equivalences on dom (${\text{\rm dom\,}}(A_{N})$ in [1, Theorem 6.10]. As $\Omega$ is bounded it is clear that the embedding $H^{3/2}(\Omega)\hookrightarrow H^{3/2-\delta/2}(\Omega)$ is compact (see, e.g., [4, Theorem 8.12.6.IV]) and hence $(A_{N}+I)^{-1}:L^{2}(\Omega)\rightarrow H^{3/2-\delta/2}(\Omega)$ is compact. Together with (3.11) we obtain that $V(A_{N}+I)^{-1}:L^{2}(\Omega)\rightarrow L^{2}(\Omega)$ is compact, that is, $V$ is relatively compact with respect to $A_{N}$. ∎
For completeness we also discuss the one-dimensional case for a finite interval $\Omega=(a,b)$. In this context we recall that the self-adjoint Neumann Laplacian in $L^{2}((a,b))$ is given by
$A_{N}f=-f^{\prime\prime},\quad{\text{\rm dom\,}}(A_{N})=\bigl{\{}f\in H^{2}((a,b))\,\big{|}\,f^{\prime}(a)=f^{\prime}(b)=0\bigr{\}}.$ (3.12)
Corollary 3.2.
Let $(a,b)$ be a finite interval, let $A_{N}$ be the self-adjoint Neumann Laplacian in $L^{2}((a,b))$, and assume that $V\not=0$ is a nonnegative function such that $V\in L^{p}((a,b))$ with $p\geq 2$. Then
$(A_{N}+\alpha V)f=-f^{\prime\prime}+\alpha Vf,\quad{\text{\rm dom\,}}(A_{N}+\alpha V)={\text{\rm dom\,}}(A_{N}),$ (3.13)
is self-adjoint in $L^{2}((a,b))$,
$V(A_{N}-zI)^{-1}\in{\mathcal{B}}_{\infty}\big{(}L^{2}((a,b))\big{)},\quad z\in\rho(A_{N}),$ (3.14)
and
$\sigma(A_{N}+\alpha V)\cap(-\infty,0)\not=\emptyset\,\text{ for any $\alpha<0$.}$ (3.15)
Proof.
Consider the densely defined closed symmetric operator
$Sf=-f^{\prime\prime},\quad{\text{\rm dom\,}}(S)=\bigl{\{}f\in H^{2}((a,b))\,\big{|}\,f(a)=f(b)=f^{\prime}(a)=f^{\prime}(b)=0\bigr{\}},$ (3.16)
[…]
$\|g\|_{L^{\infty}((a,b))}\leq C\|g\|_{H^{1}((a,b))},\quad g\in H^{1}((a,b)),$ (3.17)
one has
$\begin{split}\|Vg\|_{L^{2}((a,b))}\leq\|V\|_{L^{2}((a,b))}\|g\|_{L^{\infty}((a,b))}\leq C\|V\|_{L^{2}((a,b))}\|g\|_{H^{1}((a,b))},&\\ g\in H^{1}((a,b)),&\end{split}$ (3.18)
and hence $V:H^{1}((a,b))\rightarrow L^{2}((a,b))$ is bounded. Therefore, as $(A_{N}+I)^{-1}:L^{2}((a,b))\rightarrow H^{2}((a,b))$ is bounded and the embedding $H^{2}((a,b))\hookrightarrow H^{1}((a,b))$ is compact we see that $(A_{N}+I)^{-1}:L^{2}((a,b))\rightarrow H^{1}((a,b))$ is compact and thus also $V(A_{N}+I)^{-1}:L^{2}((a,b))\rightarrow L^{2}((a,b))$ is compact. ∎
Remark 3.3.
The observations in Corollaries 3.1 and 3.2 remain valid for more general classes of self-adjoint Laplacians. More precisely, if $\alpha\in L^{\infty}(\partial\Omega)$ is real-valued, then the Robin Laplacian
$A_{\alpha}f=-\Delta f,\quad{\text{\rm dom\,}}(A_{\alpha})=\bigl{\{}f\in H^{3/2}_{\Delta}(\Omega)\,\big{|}\,\tau_{N}f=\alpha\tau_{D}f\bigr{\}},$ (3.19)
is self-adjoint in $L^{2}(\Omega)$ and if, in addition, $A_{\alpha}$ is nonnegative and $\ker(A_{\alpha})\not=\{0\}$, then
$\sigma(A_{\alpha}+\alpha V)\cap(-\infty,0)\not=\emptyset\,\text{ for any $\alpha<0$}$ (3.20)
by Theorem 2.2 under the same integrability assumptions on $V$ as in Corollary 3.1 if $\ker(A_{\alpha})\not\subseteq\ker(V)$ holds. The latter condition is satisfied, for instance, if $V(x)>0$ for a.e. $x\in\Omega$. Similarly, in the case of a finite interval the Neumann realization $A_{N}$ of $-d^{2}/dx^{2}$ in Corollary 3.2 can be replaced by any nonnegative self-adjoint realization $A$ of $-d^{2}/dx^{2}$ in $L^{2}((a,b))$ such that $\ker(A)\not=\{0\}$. As $\ker(A)\subseteq\textrm{lin.span}\{1,x\}$ in this case, it is clear that $\ker(A)\not\subseteq\ker(V)$ holds. $\diamond$
Next, we consider the case of the Neumann (or Kirchhoff) Laplacian on a compact finite (not necessarily connected) graph $\Gamma$, which consists of $e<\infty$ edges (finite intervals) $\mathcal{E}_{n}$, $n=1,\dots,e$, and $v<\infty$ vertices $\mathcal{V}_{m}$, $m=1,\dots,v$. One recalls from [3, 24] that the self-adjoint Neumann Laplacian in $L^{2}(\Gamma)=\oplus_{n=1}^{e}L^{2}(\mathcal{E}_{n})$ is given by
$\begin{split}A_{N}f&=(-f_{n}^{\prime\prime})_{n=1}^{e},\\ {\text{\rm dom\,}}(A_{N})&=\left\{f=(f_{n})_{n=1}^{e}\,\bigg{|}\,\begin{matrix}f_{n}\in H^{2}(\mathcal{E}_{n}),\,f(x_{i})=f(x_{j}),x_{i},x_{j}\in\mathcal{V}_{m},\\ \sum_{x_{j}\in\mathcal{V}_{m}}\partial f(x_{j})=0,m=1,\dots,v,\end{matrix}\right\},\end{split}$ (3.21)
and that the multiplicity of $0\in\sigma_{p}(A_{N})$ equals the number of connected components of the metric graph $\Gamma$.
Corollary 3.4.
Let $\Gamma$ be a compact finite graph, let $A_{N}$ be the self-adjoint Neumann Laplacian in $L^{2}(\Gamma)$, and assume that $V\not=0$ is a nonnegative function such that $V\in L^{p}(\Gamma)$ with $p\geq 2$. Then
$(A_{N}+\alpha V)f=A_{N}f+\alpha Vf,\quad{\text{\rm dom\,}}(A_{N}+\alpha V)={\text{\rm dom\,}}(A_{N}),$ (3.22)
is self-adjoint in $L^{2}(\Gamma)$,
$V(A_{N}-zI)^{-1}\in{\mathcal{B}}_{\infty}\big{(}L^{2}(\Gamma)\big{)},\quad z\in\rho(A_{N}),$ (3.23)
and
$\sigma(A_{N}+\alpha V)\cap(-\infty,0)\not=\emptyset\,\text{ for any $\alpha<0$.}$ (3.24)
The proof of Corollary 3.4 is similar to that of Corollary 3.2 and hence is not repeated here.
In the next corollary we consider a perturbed Neumann Laplacian in $L^{2}((0,\infty))$, where $0\in\sigma_{p}(A)$ is an embedded eigenvalue.
Corollary 3.5.
Let
$A_{N}f=-f^{\prime\prime}+qf,\quad{\text{\rm dom\,}}(A_{N})=\bigl{\{}f\in H^{2}((0,\infty))\,\big{|}\,f^{\prime}(0)=0\bigr{\}},$ (3.25)
where
$q(x)=-\frac{2}{x^{2}+1}+\frac{8x^{2}}{(x^{2}+1)^{2}},\quad x\geq 0,$ (3.26)
and assume that $V\not=0$ is a nonnegative function such that $V\in L^{2}((0,\infty))$. Then
$(A_{N}+\alpha V)f=-f^{\prime\prime}+qf+\alpha Vf,\quad{\text{\rm dom\,}}(A_{N}+\alpha V)={\text{\rm dom\,}}(A_{N}),$ (3.27)
is self-adjoint in $L^{2}((0,\infty))$,
$V(A_{N}-zI)^{-1}\in{\mathcal{B}}_{\infty}\big{(}L^{2}((0,\infty))\big{)},\quad z\in\rho(A_{N}),$ (3.28)
$\sigma_{ess}(A_{N}+\alpha V)=\sigma_{ess}(A_{N})=[0,\infty),$ (3.29)
and
$\sigma(A_{N}+\alpha V)\cap(-\infty,0)\not=\emptyset\,\text{ for any $\alpha<0$.}$ (3.30)
Proof.
Since $q\in L^{\infty}((0,\infty))$, $A_{N}$ is self-adjoint in $L^{2}((0,\infty))$ as the same is true for the unperturbed Neumann operator $A_{0}f=-f^{\prime\prime}$, dom (${\text{\rm dom\,}}(A_{0})={\text{\rm dom\,}}(A_{N})$. It is also clear that $\infty$ is in the limit point case for the differential expression $-(d^{2}/dx^{2})+q(x)$, $x\in[0,\infty)$, and since $q\in L^{1}((0,\infty))$ it follows from [2, Proposition 6.13.7] that $\sigma_{ess}(A_{N})=\sigma_{ess}(A_{0})=[0,\infty)$. Alternatively, one can argue that the resolvent difference of the full-line Schrödinger operator associated with $-(d^{2}/dx^{2})+q(x)$, $x\in{\mathbb{R}}$, in $L^{2}({\mathbb{R}})$ and the direct sum of the corresponding two half-line Neumann operators in $L^{2}((-\infty,0))\oplus L^{2}((0,\infty))$ is a rank-one operator and combine this with the fact that $q(x)=q(-x)$, $x\in[0,\infty)$, and the full-line Schrödinger operator has essential spectrum equal to $[0,\infty)$ as $\lim_{x\to\pm\infty}q(x)=0$. Moreover, it is easy to see that 00 is a simple eigenvalue of $A_{N}$ with corresponding eigenfunction
$k(x)=\frac{1}{x^{2}+1},\quad x\in[0,\infty),$ (3.31)
and it follows from
$A_{N}=B^{*}B\geq 0,$ (3.32)
that $A_{N}$ is nonnegative. Here,
$\begin{split}&Bf=f^{\prime}+\phi f,\quad f\in{\text{\rm dom\,}}(B)=H^{1}([0,\infty)),\\ &B^{*}g=-g^{\prime}+\phi g,\quad g\in{\text{\rm dom\,}}(B^{*})=H^{1}_{0}([0,\infty)),\end{split}$ (3.33)
where
$\phi(x)=k^{\prime}(x)/k(x)=-\frac{2x}{x^{2}+1},\quad x\in[0,\infty).$ (3.34)
We also note that the condition $\ker(A_{N})\not\subseteq\ker(V)$ in Theorem 2.2 is satisfied for the multiplication operator $V$ as otherwise $V=0$. We claim that $V$ is relatively compact with respect to $A_{N}$. In fact, for $z\in{\mathbb{C}}\setminus[0,\infty)$ we have the identity
$(A_{N}-zI)^{-1}=(A_{0}-zI)^{-1}-(A_{0}-zI)^{-1}q(A_{N}-zI)^{-1}$ (3.35)
and $V(A_{0}-zI)^{-1}\in{\mathcal{B}}_{\infty}(L^{2}((0,\infty)))$ by [29, Problem 41] (for the half-line), and thus also $V(A_{N}-zI)^{-1}\in{\mathcal{B}}_{\infty}(L^{2}((0,\infty)))$. This implies $\sigma_{ess}(A_{N}+\alpha V)=\sigma_{ess}(A_{N})=[0,\infty)$ and hence (3.30) follows from Theorem 2.2. ∎
Acknowledgments. We are indebted to Petr Siegl for fruitful discussions and helpful remarks. J.B. gratefully acknowledges financial support by the Austrian Science Fund (FWF): P33568-N. J.B. is also most grateful for a stimulating research stay at Baylor University, where some parts of this paper were written in October of 2023. F.G. and H.S. gratefully acknowledge kind invitations to the Institute for Applied Mathematics at the Graz University of Technology, Austria.